Find the solution by substitution:
3x+8y=22
10x-2y=16

Answer by 24HoursTutor.com(40)

3x+8y=22 and
10x-2y=16
Let us find the value of y in terms of x from the second equation so that we can substitute it in the first one.
10x-2y=16
10x-16=2y{divide everything by 2 now}
5x-8=y
y=5x-8
Now, substitute this value of y in the first equation :
3x + 8y = 22
3x + 8(5x-8) = 22
3x + 40x-64 = 22
43x - 64 = 22
43x =22+64
43x =86
x = 86/43 = 2
x=2
Now substitute the value of x in the following equation to get y.
y=5x-8
y=5(2)-8
y=10-8=2
Ans. Both the values of x and y are 2.
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr